OpenStudy (anonymous):

What is the standard form of the equation of the conic given by 2x^2 + 2y^2 + 4x - 12y – 22 = 0?

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

What's next?

OpenStudy (aum):

2x^2 + 2y^2 + 4x - 12y – 22 = 0 divide throughout by 2: x^2 + y^2 + 2x - 6y - 11 = 0 (x^2 + 2x) + (y^2 - 6y) – 11 = 0 complete the square of (x^2 + 2x) complete the square of (y^2 - 6y) First complete the two squares and we can go to the next step.

OpenStudy (anonymous):

How do I complete the squares?

OpenStudy (aum):

Look under "steps for completing the square" here: http://www.regentsprep.org/regents/math/algtrig/ate12/completesqlesson.htm

OpenStudy (anonymous):

Is it (x+1)^2-1 and (y-3)^2-9

OpenStudy (aum):

correct.

OpenStudy (anonymous):

Awesome, so what's next?

OpenStudy (aum):

(x^2 + 2x) + (y^2 - 6y) – 11 = 0 (x + 1)^2 - 1 + (y - 3)^2 - 9 - 11 = 0 (x+1)^2 + (y-3)^2 = 21 \((x+1)^2 + (y-3)^2 = (\sqrt{21})^2 \) Compare it to: \((x-h)^2 + (y-k)^2 = r^2 \) This is a circle with center at (-1, 3) and radius \(\sqrt{21}\)

OpenStudy (aum):

h = -1, k = 3, r = \(\sqrt{21}\) where (h, k) is the center of the circle and r is its radius.

OpenStudy (aum):

They are just asking you to put the equation in standard form and so the answer is: \((x+1)^2 + (y-3)^2 = (\sqrt{21})^2\) They don't ask you to identify the equation. So the rest are all FYI.

OpenStudy (anonymous):

Can you help with one more?